OpenStudy (anonymous):

show steps used to solve 5n + 6 = -4

OpenStudy (anonymous):

5n+6=-4 5n +6 -6 = -4-6 whatever you do to one side of the = you must do to the other 5n = -10 5/5n = - 10/5 whatever you do to one side of the = you must do to the other 5 divided into 5 is 1 or just n n= -10/5 5 divided into 10 is 2........don't forget the minus sign n=-2

OpenStudy (thesecret20111):

Thanks for fixing my answer! I somehow forgot the minus :S